# TexHOI: Reconstructing Textures of 3D Unknown Objects in Monocular Hand-Object Interaction Scenes

### v1

@@ Abstract
Reconstructing 3D models of dynamic, real-world objects with high-fidelity textures from monocular frame sequences has been a challenging problem in recent years. This difficulty stems from factors such as shadows, indirect illumination, and inaccurate object-pose estimations due to occluding hand-object interactions. To address these challenges, we propose a novel approach that predicts the hand’s impact on environmental visibility and indirect illumination on the object’s surface albedo. Our method first learns the geometry and low-fidelity texture of the object, hand, and background through composite rendering of radiance fields. Simultaneously, we optimize the hand and object poses to achieve accurate object-pose estimations. We then refine physics-based rendering parameters—including roughness, specularity, albedo, hand visibility, skin color reflections, and environmental illumination—to produce precise albedo, and accurate hand illumination and shadow regions. Our approach surpasses state-of-the-art methods in texture reconstruction and, to the best of our knowledge, is the first to account for hand-object interactions in object texture reconstruction. Please check our source code at: https://anonymous.4open.science/r/TexHOI-BCCE
Keywords: Radiance Field, Differentiable Renderer, Spherical Gaussian, Albedo Prediction
@@ 1 Introduction
In recent years, there has been an increase in the availability of standardized datasets designed to capture hand-object interactions, such as HO3D [hampali2020honnotate] and DexYCB [chao2021dexycb], which are built on YCB-models [xiang2017posecnn], or Contact-Pose [brahmbhatt2020contactpose] dataset. These datasets provide crucial insights into how hands manipulate everyday objects, like sugar boxes and snack containers, in dynamic scenarios. However, a major challenge remains in effectively utilizing these datasets due to the complex shadowing and reflection effects that hands cast on object surfaces. Accurately capturing and rendering these effects is critical for a wide range of applications, from enhancing realism in virtual and augmented reality environments to improving robotic manipulation and object recognition systems. Several recent works have focused on reconstructing hand-object interactions. Methods like BundleSDF [wen2023bundlesdf], DiffHOI [ye2023diffusion], and HOLD [fan2024hold] primarily address geometry reconstruction. However, these methods struggle with generating accurate textures under varying environmental illumination conditions. Moreover, shadows and indirect illumination caused by hand movements introduce significant errors in texture prediction. To the best of our knowledge, no recent work explicitly addresses the influence of shadows cast by dynamic elements like hands on 3D object reconstruction, particularly texture reconstruction. Other works such as PhySG [zhang2021physg], InvRender [zhang2022modeling], and RefNeRF [refnerf] focus on modeling the influence of environmental illumination on object surfaces based on material properties, typically represented as Spherical Harmonics or Spherical Gaussian. However, these methods do not account for interacting non-rigid parts like hands, making them unable to handle hand-induced shadows and reflections. Additionally, they rely on precise camera pose estimations based on object-coordinate systems, which can introduce errors akin to motion blur when these assumptions are violated. Given the above challenges, our primary research question is: How can we achieve accurate texture and geometry predictions in dynamic hand-object interaction scenarios using monocular camera data?
@@
Our approach begins by optimizing both hand and object poses using a composite rendering technique that separately models the hand, object, and background. In this process, we simultaneously refine hand and object poses while learning neural radiance fields for each entity, ensuring accurate geometric and pose representations. Using these fine-tuned poses, we represent the environmental illumination with Spherical Gaussian (SG) representations. This allows us to model the interaction between the object’s material properties — such as specularity and roughness — and the environment. To efficiently compute the shadows and reflections caused by interacting hands, we introduce a novel representation of the hand using 108 parameterizable spheres. For each surface point on the object, we calculate the visibility and indirect illumination caused by each sphere by measuring the fraction of its projection that occludes the SG on the object’s surface. This method avoids the need for time-consuming ray-tracing [zhang2021nerfactor] or learning pseudo-ground truth visibility using MLPs [zhang2022modeling]. We then combine all the elements in the physics-based rendering (PBR) framework to obtain the final object color, which is compared with the ground truth camera-captured images to optimize all those physics-based rendering parameters, i.e., roughness, specularity, albedo, hand visibility, skin color reflections, and environmental illumination.
@@ 2 Related Works
Creating realistic rendered images of a target object, for its re-usability in varying virtual environment has been a major focus in the study of computer graphics for much of its history. Inverse rendering of an object from their visual data refers to predicting camera poses and scene descriptions, including the geometry of the object, the reflectance of the object surface, which consists of its albedo texture and other microfacet surface properties like roughness, specularity, etc., and environment lighting used to illuminate the scene [marschner1998inverse]. Many research works have built the basic framework for upcoming advancements in this area of research.
@@ 2.1 Traditional Inverse Rendering for 3D Object Reconstruction
Traditional inverse rendering approaches broadly divided the research field into three categories - inverse lighting to predict the environment light that illuminates the scene in visual input such as images, inverse reflectometry to predict object texture or object BRDF from visual input, or shape reconstruction to obtain the geometry of the object. Traditional approaches only focus on solving one of the three mentioned problems. They may solve inverse lighting problem [kawai1993radioptimization] which allows re-lighting of rendered scenes under different environment illumination [nimeroff1995efficient, teo1997cient]. Some other works solve inverse reflectometry problems to obtain realistic texture [turk1991generating, witkin1991reaction] or accurate reflective model [curless1996volumetric, eck1995multiresolution] on the object surface. Meanwhile, other approaches optimize object geometry [ozyecsil2017survey, su2014estimating, james2012straightforward, westoby2012structure]. These approaches assume other aspects of the inverse rendering, except their focus problem, to be known. To our knowledge, no traditional approach comprehensively addresses all aspects of the inverse rendering problem in a single study. Recent advancements in neural representations address the comprehensive computation of these different aspects.
@@ 2.2 Neural Representation in Inverse Rendering
In recent years, several recent neural network-based research works have sparked a wave of innovative methodologies to solve inverse rendering problems. These include volumetric rendering based approaches like NeRF [mildenhall2021nerf], NeuS [wang2021neus], DVGO [sun2022improved], TensoRF [chen2022tensorf], Instant-NGP [muller2022instant], etc., surface rendering based approaches like IDR [yariv2020multiview] or point-based inverse rendering approaches like 3D-Gaussian-Splatting [kerbl20233d]. Point-based inverse rendering is more efficient and helps in real-time optimization, but suffers from high-detail geometry and texture prediction. Moreover, these approaches give a general framework solution for inverse rendering problems but do not solve complex problems that arise from varying material properties of objects, environment illumination in the scene, or occlusion and visibility handling in multi-object interactions in the scene. Applying the physics-based knowledge, and addressing the composition of multiple-object interactions in the scene further improves this research area, as will be seen in the next two subsections.
@@ 2.3 Physics-based Inverse Rendering
where $\omega_{i}$ is a direction in hemi-sphere $\Omega^{+}(n)$, defined with the $up$ direction as surface normal $n$. $L(\omega_{i},x)$ is obtained by inverse lighting, $\phi(x;\omega_{o},\omega_{i})$ is obtained by inverse reflectometry and normal $n$ is obtained by shape reconstruction. Occlusions between the illumination source and surface point $x$ may break down $L(x)$ term, requiring visibility and indirect illumination calculation. Many recent approaches have done extensive research on improving the inverse rendering output by predicting the above terms of the PBR equation. NeRV [srinivasan2021nerv] assumes known illumination. Other methods like NeRD [boss2021nerd], Neural Ray-Tracing [knodt2021neural], NeRFactor [zhang2021nerfactor], Neural-PIL [boss2021neural], PS-NeRF [yang2022ps], Neilf [yao2022neilf], SAMURAI [boss2022samurai], NVDiffRec [hasselgren2022shape], Neural-PBIR [sun2023neural], TensoIR [jin2023tensoir], TensoSDF [li2024tensosdf], MIRReS [dai2024mirres], IRON [zhang2022iron] etc. use volumetric-rendering based methods to optimize neural radiance fields, and extract geometry, illumination, texture and reflectance model. NeRO [liu2023nero], NeP [wang2024inverse], PBIR-NIE [cai2024pbir] extend these methods to glossy or reflective surfaces. Relightable-3D-Gaussian [gao2023relightable] uses 3D Gaussian splats, and methods like IntrinsicAnything [chen2024intrinsicanything] learn diffusion priors to obtain accurate albedo and material properties. Methods like PhySG [zhang2021physg] and InvRender [zhang2022modeling] use surface-based rendering approaches and approximate illumination and BRDF as Spherical Gaussians (SGs). An SG comprises of a lobe with its direction $\xi$, intensity $\mu$ and width $\lambda$, and is defined on a spherical surface as follows:
$G(v;\xi,\mu,\lambda)=\mu e^{\lambda(v\cdot\xi-1)},$ (2)
where, $v$ is any direction on the sphere and $G(v)$ is the value on the sphere along direction $v$. The above methods either ignore visibility computation [zhang2021physg], perform computationally-extensive Monte-Carlo ray tracing [zhang2021nerfactor], or estimate pseudo-ground truth visibility using a Multi-layer Perceptron (MLP) [zhang2022modeling]. Monte-Carlo ray-tracing can take up to  30 minutes per frame on a single GPU to train. L-Tracing [chen2022tracing] utilizes sphere-tracing, and the concavity/convexity of geometry to avoid learning visibility. However, both InvRender [zhang2022modeling] and L-Tracing [chen2022tracing] work for a single static rigid object, and fail to handle multiple dynamic object visibility estimation. NeRF-DS [yan2023nerf] handles dynamic objects but does not utilize the PBR equation to obtain the reflectometry information in the scene. Instead, the method is only restricted to glossy objects. Moreover, all the above methods assume a single rigid object with accurate pose estimations in the scene. These methods do not handle multiple dynamic object interactions in the scene.
@@ 2.4 Handling Dynamic Elements in Multi-Object Interaction
To handle multiple object interactions, along with background, many recent researches have come up, with compositional-NeRF methods [wang2023learning, driess2023learning, wu2022object, yang2021learning]. HandNeRF [guo2023handnerf] extends this compositional-NeRF methodology for multiple interacting hands. Other methods named HandNeRF [choi2024handnerf], BundleSDF [wen2023bundlesdf], HOLD [fan2024hold], etc. focus on dynamic hand-object reconstruction. However, their primary focus is geometry reconstruction, and do not focus on high-fidelity texture reconstruction. Hence, they do not focus on physics-based accuracy in results. Our research primarily focuses on 3D geometry and high-fidelity texture reconstruction of hand-object interactions from single-camera data capture sequences. We aim to improve the output by resolving the baking of hand effects into the rendered object texture, and obtain a high-fidelity object albedo prediction following a two-stage training process. In the first stage, we fine-tune hand and object poses using a compositional volumetric rendering approach [fan2024hold]. In the second stage, we use a surface rendering approach with Spherical Gaussians to obtain different elements in the PBR equation, including hand effects on the object surface. In Sec. 4, we will describe the above two stages of training in-depth.
@@ 3.2 Surface Rendering
The surface rendering focuses on identifying the exact intersection point $\hat{t}$ of the ray $r(\hat{t})=o+\hat{t}d$ with the implicit surface geometry using the sphere tracing approach. Once the intersection point is determined, the color at that point is predicted as $c(r(\hat{t}),d)$. After calculating the surface color, it is assigned to the corresponding pixel in the image, forming the final rendered image. Methods like IDR [yariv2020multiview], PhySG [zhang2021physg], etc. use surface rendering.
@@ 4.1 Stage 1: Compositional Neural Radiance Field for Pose Refinement
This stage utilizes three Neural Radiance Fields (NeRF). NeRF for hand optimizes the hand pose transformation and generates hand from monocular camera frames in MANO [romero2022embodied] canonical pose. Similarly, a NeRF for the object optimizes the object pose transformation and generates the rigid object in its canonical pose, or world origin. Meanwhile, another NeRF learns background image representation. Finally, the outputs of the three NeRFs are composed to generate a composite rendered image similar to the ground-truth image frame. In this section, we will discuss in detail the above-mentioned NeRFs and their composition.
@@ 4.1.2 Hand NeRF
Calculating the radiance field for hand NeRF follows a similar setup as object NeRF. The points sampled along the camera ray are inverse-transformed to the hand’s canonical pose, i.e., stretched hand, using MANO [romero2022embodied] pose parameters. Hence, we again need to perform the inverse transformation for hand. Given transformation matrices $T_{i}$ for $i^{th}$ of $n_{jnt}$ joints, with interpolated skinning weight $w_{i}(x^{\prime})$ on the observation space point $x^{\prime}$, its corresponding canonical space point $x_{hand}$ is calculated as follows:
$x_{hand}=(\Sigma_{i=1}^{{\color[rgb]{0,0,1}n_{jnt}}}(w_{i}(x^{\prime})T_{i})^{-1}x^{\prime}.$ (5)
@@ 4.1.3 Compositional NeRF
Mathematically, sampling $n$ points for object and inverse transforming the points into object-coordinate system 4.1.1, and $n$ points for hand and performing linear inverse blend-skinning into hand’s canonical coordinate system 4.1.2, we obtain the foreground (hand and object) color $C_{fg}$ and foreground mask $M_{fg}$. The values are calculated by sorting the $2n$ points, and the predicted density $\sigma_{i}$ and radiance $col_{i}$ values of each point. Similarly, using the calculated foreground mask $M_{fg}$ and the background color $C_{bg}$, we obtain final predicted color along a ray $r$ as follows:
@@
$C_{fg}(r)=\Sigma_{i=1}^{2n}\tau_{i}{\color[rgb]{0,0,1}col_{i}}.$ (7)
Here $\tau_{i}=exp(-\Sigma_{j<i}\sigma_{j}\delta_{j})(1-exp(-\sigma_{i}\delta_{i}))$ and $\delta_{i}$ is the distance between $i^{th}$ and $(i+1)^{th}$ sample points. Similarly, the foreground mask is given as:

### v2

@@ Abstract
Reconstructing 3D models of dynamic, real-world objects with high-fidelity textures from monocular frame sequences has been a challenging problem in recent years. This difficulty stems from factors such as shadows, indirect illumination, and inaccurate object-pose estimations due to occluding hand-object interactions. To address these challenges, we propose a novel approach that predicts the hand’s impact on environmental visibility and indirect illumination on the object’s surface albedo. Our method first learns the geometry and low-fidelity texture of the object, hand, and background through composite rendering of radiance fields. Simultaneously, we optimize the hand and object poses to achieve accurate object-pose estimations. We then refine physics-based rendering parameters—including roughness, specularity, albedo, hand visibility, skin color reflections, and environmental illumination—to produce precise albedo, and accurate hand illumination and shadow regions. Our approach surpasses state-of-the-art methods in texture reconstruction and, to the best of our knowledge, is the first to account for hand-object interactions in object texture reconstruction. Please check our work at: https://alakhag.github.io/TexHOI-website/
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
Index Terms:
Radiance Field, Differentiable Renderer, Spherical Gaussian, Albedo Prediction
@@ 1 Introduction
In recent years, there has been an increase in the availability of standardized datasets designed to capture hand-object interactions, such as HO3D [21] and DexYCB [7], which are built on YCB-models [57], or Contact-Pose [5] dataset. These datasets provide crucial insights into how hands manipulate everyday objects, like sugar boxes and snack containers, in dynamic scenarios. However, a major challenge remains in effectively utilizing these datasets due to the complex shadowing and reflection effects that hands cast on object surfaces. Accurately capturing and rendering these effects is critical for a wide range of applications, from enhancing realism in virtual and augmented reality environments to improving robotic manipulation and object recognition systems. Several recent works have focused on reconstructing hand-object interactions. Methods like BundleSDF [53], DiffHOI [65], and HOLD [17] primarily address geometry reconstruction. However, these methods struggle with generating accurate textures under varying environmental illumination conditions. Moreover, shadows and indirect illumination caused by hand movements introduce significant errors in texture prediction. To the best of our knowledge, no recent work explicitly addresses the influence of shadows cast by dynamic elements like hands on 3D object reconstruction, particularly texture reconstruction. Other works such as PhySG [67], InvRender [69], and RefNeRF [48] focus on modeling the influence of environmental illumination on object surfaces based on material properties, typically represented as Spherical Harmonics or Spherical Gaussian. However, these methods do not account for interacting non-rigid parts like hands, making them unable to handle hand-induced shadows and reflections. Additionally, they rely on precise camera pose estimations based on object-coordinate systems, which can introduce errors akin to motion blur when these assumptions are violated. Given the above challenges, our primary research question is: How can we achieve accurate texture and geometry predictions in dynamic hand-object interaction scenarios using monocular camera data?
@@
Our approach begins by optimizing both hand and object poses using a composite rendering technique that separately models the hand, object, and background. In this process, we simultaneously refine hand and object poses while learning neural radiance fields for each entity, ensuring accurate geometric and pose representations. Using these fine-tuned poses, we represent the environmental illumination with Spherical Gaussian (SG) representations. This allows us to model the interaction between the object’s material properties — such as specularity and roughness — and the environment. To efficiently compute the shadows and reflections caused by interacting hands, we introduce a novel representation of the hand using 108 parameterizable spheres. For each surface point on the object, we calculate the visibility and indirect illumination caused by each sphere by measuring the fraction of its projection that occludes the SG on the object’s surface. This method avoids the need for time-consuming ray-tracing [68] or learning pseudo-ground truth visibility using MLPs [69]. We then combine all the elements in the physics-based rendering (PBR) framework to obtain the final object color, which is compared with the ground truth camera-captured images to optimize all those physics-based rendering parameters, i.e., roughness, specularity, albedo, hand visibility, skin color reflections, and environmental illumination.
@@ 2 Related Works
Creating realistic rendered images of a target object, for its re-usability in varying virtual environment has been a major focus in the study of computer graphics for much of its history. Inverse rendering of an object from their visual data refers to predicting camera poses and scene descriptions, including the geometry of the object, the reflectance of the object surface, which consists of its albedo texture and other microfacet surface properties like roughness, specularity, etc., and environment lighting used to illuminate the scene [34]. Many research works have built the basic framework for upcoming advancements in this area of research.
@@ 2.1 Traditional Inverse Rendering for 3D Object Reconstruction
Traditional inverse rendering approaches broadly divided the research field into three categories - inverse lighting to predict the environment light that illuminates the scene in visual input such as images, inverse reflectometry to predict object texture or object BRDF from visual input, or shape reconstruction to obtain the geometry of the object. Traditional approaches only focus on solving one of the three mentioned problems. They may solve inverse lighting problem [26] which allows re-lighting of rendered scenes under different environment illumination [37, 46]. Some other works solve inverse reflectometry problems to obtain realistic texture [47, 55] or accurate reflective model [13, 16] on the object surface. Meanwhile, other approaches optimize object geometry [38, 43, 24, 54]. These approaches assume other aspects of the inverse rendering, except their focus problem, to be known. To our knowledge, no traditional approach comprehensively addresses all aspects of the inverse rendering problem in a single study. Recent advancements in neural representations address the comprehensive computation of these different aspects.
@@ 2.2 Neural Representation in Inverse Rendering
In recent years, several recent neural network-based research works have sparked a wave of innovative methodologies to solve inverse rendering problems. These include volumetric rendering based approaches like NeRF [35], NeuS [51], DVGO [45], TensoRF [8], Instant-NGP [36], etc., surface rendering based approaches like IDR [64] or point-based inverse rendering approaches like 3D-Gaussian-Splatting [27]. Moreover, these approaches give a general framework solution for inverse rendering problems but do not solve complex problems that arise from varying material properties of objects, environment illumination in the scene, or occlusion and visibility handling in multi-object interactions in the scene. Applying the physics-based knowledge, and addressing the composition of multiple-object interactions in the scene further improves this research area, as will be seen in the next two subsections.
@@ 2.3 Physics-based Inverse Rendering
where $\omega_{i}$ is a direction in hemi-sphere $\Omega^{+}(n)$, defined with the $up$ direction as surface normal $n$. $L(\omega_{i},x)$ is obtained by inverse lighting, $\phi(x;\omega_{o},\omega_{i})$ is obtained by inverse reflectometry and normal $n$ is obtained by shape reconstruction. Occlusions between the illumination source and surface point $x$ may break down $L(x)$ term, requiring visibility and indirect illumination calculation. Many recent approaches have done extensive research on improving the inverse rendering output by predicting the above terms of the PBR equation. NeRV [42] assumes known illumination. Other methods like NeRD [2], Neural Ray-Tracing [29], NeRFactor [68], Neural-PIL [4], PS-NeRF [60], Neilf [63], SAMURAI [3], NVDiffRec [22], Neural-PBIR [44], TensoIR [25], TensoSDF [30], MIRReS [14], IRON [66] etc. use volumetric-rendering based methods to optimize neural radiance fields, and extract geometry, illumination, texture and reflectance model. NeRO [33], NeP [49], PBIR-NIE [6] extend these methods to glossy or reflective surfaces. Relightable-3D-Gaussian [18] uses 3D Gaussian splats, and methods like IntrinsicAnything [9] learn diffusion priors to obtain accurate albedo and material properties. Methods like PhySG [67] and InvRender [69] use surface-based rendering approaches and approximate illumination and BRDF as Spherical Gaussians (SGs). An SG comprises of a lobe with its direction $\xi$, intensity $\mu$ and width $\eta$, and is defined on a spherical surface as follows:
$G(v;\xi,\mu,{\eta})=\mu e^{{\eta}(v\cdot\xi-1)},$ (2)
where, $v$ is any direction on the sphere and $G(v)$ is the value on the sphere along direction $v$. The above methods either ignore visibility computation [67], perform computationally-extensive Monte-Carlo ray tracing [68], or estimate pseudo-ground truth visibility using a Multi-layer Perceptron (MLP) [69]. Monte-Carlo ray-tracing can take up to  30 minutes per frame on a single GPU to train. L-Tracing [10] utilizes sphere-tracing, and the concavity/convexity of geometry to avoid learning visibility. However, both InvRender [69] and L-Tracing [10] work for a single static rigid object, and fail to handle multiple dynamic object visibility estimation. NeRF-DS [58] handles dynamic objects but does not utilize the PBR equation to obtain the reflectometry information in the scene. Instead, the method is only restricted to glossy objects. Moreover, all the above methods assume a single rigid object with accurate pose estimations in the scene. These methods do not handle multiple dynamic object interactions in the scene.
@@ 2.4 Handling Dynamic Elements in Multi-Object Interaction
To handle multiple object interactions, along with background, many recent researches have come up, with compositional-NeRF methods [52, 15, 56, 59]. HandNeRF [20] extends this compositional-NeRF methodology for multiple interacting hands. Other methods named HandNeRF [12], BundleSDF [53], HOLD [17], etc. focus on dynamic hand-object reconstruction. However, their primary focus is geometry reconstruction, and do not focus on high-fidelity texture reconstruction. Hence, they do not focus on physics-based accuracy in results. Our research primarily focuses on 3D geometry and high-fidelity texture reconstruction of hand-object interactions from single-camera data capture sequences. We aim to improve the output by resolving the baking of hand effects into the rendered object texture, and obtain a high-fidelity object albedo prediction following a two-stage training process. In the first stage, we fine-tune hand and object poses using a compositional volumetric rendering approach [17]. In the second stage, we use a surface rendering approach with Spherical Gaussians to obtain different elements in the PBR equation, including hand effects on the object surface. In Sec. 4, we will describe the above two stages of training in-depth.
@@ 3.2 Surface Rendering
The surface rendering focuses on identifying the exact intersection point $\hat{t}$ of the ray $r(\hat{t})=o+\hat{t}d$ with the implicit surface geometry using the sphere tracing approach. Once the intersection point is determined, the color at that point is predicted as $c(r(\hat{t}),d)$. After calculating the surface color, it is assigned to the corresponding pixel in the image, forming the final rendered image. Methods like IDR [64], PhySG [67], etc. use surface rendering.
@@ 4.1 Stage 1: Compositional Neural Radiance Field for Pose Refinement
This stage utilizes three Neural Radiance Fields (NeRF). NeRF for hand optimizes the hand pose transformation and generates hand from monocular camera frames in MANO [39] canonical pose. Similarly, a NeRF for the object optimizes the object pose transformation and generates the rigid object in its canonical pose, or world origin. Meanwhile, another NeRF learns background image representation. Finally, the outputs of the three NeRFs are composed to generate a composite rendered image similar to the ground-truth image frame. In this section, we will discuss in detail the above-mentioned NeRFs and their composition.
@@ 4.1.2 Hand NeRF
Calculating the radiance field for hand NeRF follows a similar setup as object NeRF. The points sampled along the camera ray are inverse-transformed to the hand’s canonical pose, i.e., stretched hand, using MANO [39] pose parameters. Hence, we again need to perform the inverse transformation for hand. Given transformation matrices $T_{i}$ for $i^{th}$ of $n_{jnt}$ joints, with interpolated skinning weight $w_{i}(x^{\prime})$ on the observation space point $x^{\prime}$, its corresponding canonical space point $x_{hand}$ is calculated as follows:
$x_{hand}=(\Sigma_{i=1}^{{n_{jnt}}}(w_{i}(x^{\prime})T_{i})^{-1}x^{\prime}.$ (5)
@@ 4.1.3 Compositional NeRF
Mathematically, sampling $n$ points for object and inverse transforming the points into object-coordinate system 4.1.1, and $n$ points for hand and performing linear inverse blend-skinning into hand’s canonical coordinate system 4.1.2, we obtain the foreground (hand and object) color $C_{fg}$ and foreground mask $M_{fg}$. The values are calculated by sorting the $2n$ points, and the predicted density $\sigma_{i}$ and radiance $c_{i}$ values of each point. Similarly, using the calculated foreground mask $M_{fg}$ and the background color $C_{bg}$, we obtain final predicted color along a ray $r$ as follows:
@@
$C_{fg}(r)=\Sigma_{i=1}^{2n}\tau_{i}c_{i}.$ (7)
Here $\tau_{i}=exp(-\Sigma_{j<i}\sigma_{j}\delta_{j})(1-exp(-\sigma_{i}\delta_{i}))$ and $\delta_{i}$ is the distance between $i^{th}$ and $(i+1)^{th}$ sample points. Similarly, the foreground mask, is given as:
@@
Individually, mask for hand along a ray $r$ can be obtained using Hand NeRF in Sec. 4.1.2 as $M_{hand}(r)$. Similarly, masks for object (as shown in Fig. 2(b)) along a ray $r$ can be obtained using Object NeRF in Sec. 4.1.1 as $M_{obj}(r)$. Whereas, the background mask can be obtained using the foreground mask as $M_{bg}(r)=1-M_{fg}(r)$. Finally, along a ray $r$, the one-hot encoding of hand mask $M_{hand}(r)$, object mask $M_{obj}(r)$ and background mask $M_{bg}(r)$ gives us the predicted segmentation $M(r)$:
$M(r)=\arg max\{M_{hand}(r),M_{obj}(r),M_{bg}(r)\}.$ (9)
For stage 2, we also need visible object mask by following similar methodology as Eqs. 78. We can set the color prediction of the sample points for object to $1$, and color predictions for hand as 00. Along a ray $r$, this allows only the rays from the visible object regions to transmit, and stops the transmittance from the object surface points occluded by hands, giving us visible object mask $M_{ho}$, as shown in Fig. 2(c).